# Towards Black-Box Membership Inference Attack for Diffusion Models

### v3

@@ 1 Introduction
Recently, there has been a surge in the popularity of generative models, with diffusion models in particular, gaining huge attention within the AI community [42, 45, 46]. These models have demonstrated remarkable capabilities across various tasks, including unconditional image generation [14, 44], text-to-image generation [36, 53, 31] and image-to-image generation [37]. This surge has given rise to powerful AI art models such as DALL-E 2 [35], Stable Diffusion [36], and Imagen [38]. AI-generated art holds a promising future and is expected to have a widespread impact.
Effective training of diffusion models requires high-quality data. It is thus crucial to design an algorithm that can identify whether a specific artwork has been used during the training of a model, thereby providing protection for these artworks and detecting misuse of data. This is especially important due to the rapid growth of generative models, which has raised concerns over intellectual property (IP) rights, data privacy, and the ethical implications of training on copyrighted or proprietary content without consent. As these models are increasingly deployed across industries, detecting whether a specific piece of content was used in training can help prevent unauthorized use of artistic works, protecting creators’ copyrights and ownership rights. This is a classic problem in the field of machine learning, first introduced by [41] and named “membership inference attack”.
A series of studies have been conducted on membership inference attacks against diffusion models. [16] was the first to examine this issue, utilizing the loss function values of diffusion models to determine whether an image is in the training set. [6]and [20] extended this work by relaxing the assumptions about model access requirements.
@@
    We evaluate our method using DDIM [44] and Stable Diffusion [36] models on classical datasets, including CIFAR10/100 [22], STL10-Unlabeled [2], LAION-5B [40], etc. Our method outperforms the previous methods.
@@
    We extend both existing algorithms and our own algorithm to the Diffusion Transformer [33] architecture, implementing the membership inference attack within this model framework for the first time. Experimental results demonstrate that our algorithm is consistently effective.
@@ 2 Related Works
The diffusion model, initially proposed by [42], has achieved remarkable results in producing high-quality samples across a variety of domains. This ranges from image generation [45, 46, 4], audio synthesis [34, 21, 17], and video generation [13, 15, 50], etc. Among existing diffusion models, the Denoising Diffusion Probabilistic Model (DDPM) [14] is one of the most frequently adopted. This approach introduces a dual-phase process for image generation: initially, a forward process gradually transforms training data into pure noise, followed by a reverse process that meticulously reconstructs the original data from this noise. Building on this model, there have been numerous follow-up studies, such as Stable Diffusion [36], which compresses images into a latent space and generates images based on text, and the Denoising Diffusion Implicit Models (DDIM) [44], which removes Gaussian randomness to accelerate the sampling generation process. These advancements demonstrate the versatility and potential of diffusion models.
@@
In the era of big data, preserving data privacy is paramount. The training of diffusion models may involve sensitive datasets like artists’ artworks, which are protected by copyright laws. Membership inference attacks, initially introduced by [41], serve as an effective means to detect potential misuse of data without proper authorization. Its objective is to ascertain whether a particular data sample participated in the training phase of a target model. This approach is instrumental in probing privacy breaches and identifying illicit data utilization. Researchers primarily focus on membership inference attacks for classification models [39, 52, 26, 24], embedding models [43, 7, 28], and generative models [10, 12, 1].
In the domain of membership inference attacks against diffusion models, [51, 16] use a white-box approach, which assumes access to the entire diffusion model and utilizes loss and likelihood to determine whether a sample is in the training set. [6, 20, 47] have relaxed these requirements, eliminating the need for the entire model. They leverage the insight that samples within the training set yield more accurate noise predictions, thereby achieving high accuracy in membership inference attack tasks. However, they also require the outputs of the U-net, as it is necessary to obtain the noise predictions of intermediate steps. Recently, [32] proposed a black-box membership inference attack method against diffusion models. Their method identifies whether a specific image is in a finetuning dataset of 100 images by calculating the difference between the generated image and the target image, using the corresponding prompt as input. Their approach leverages the model’s tendency to memorize finetuning images and generate similar outputs. In contrast, we focus on detecting whether an image is in the pretraining dataset, where a pre-trained model often produces diverse outputs for the same prompt, making detection more challenging.
@@ 3 Preliminary
In this section, we begin by introducing the notations used for several popular diffusion models. We first introduce the Denoising Diffusion Probabilistic Model (DDPM) [14]. Then, we extend to the Denoising Diffusion Implicit Model (DDIM) [44] and Stable Diffusion [36], which are variants of DDPM used to accelerate image generation or generate images grounded in text descriptions. Lastly, we discuss Diffusion Transformer [33], a model that replaces the U-net architecture with a transformer and achieves higher-quality image generation.
@@
$x_{t-1}=\phi_{\theta}(x_{t},t)=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x_{t},t)\,,$ (1)
@@
Stable Diffusion leverages a variational autoencoder (VAE) [19] to encode images into a latent space and perform diffusion in this compressed space. The model uses a text encoder to guide the diffusion process, enabling text-to-image generation:
@@
Diffusion Transformer leverages the Vision Transformer [5] structure to replace the U-net architecture traditionally used in diffusion models for noise prediction. Its training and sampling methods remain consistent with DDIM, with the only difference being the replacement of noise prediction network $\epsilon_{\theta}$ with $\epsilon_{\tilde{\theta}}$, where $\tilde{\theta}$ represents a Vision Transformer-based architecture. This approach further enhances the generation quality and ensured that the model possesses good scalability properties.
@@ 4.1 The variation API for Diffusion Models
Most previous works on membership inference attacks against diffusion models aim to prevent data leakage and hence rely on thresholding the model’s training loss. For instance, [16] involves a direct comparison of image losses, while [6, 20] evaluates the accuracy of the model’s noise prediction at initial or intermediate steps. However, the required access to the model’s internal U-net structure prevents applications from copyright protection because most servers typically provide only black-box API access.
@@
Definition 1 (The variation API).
We define the variation API $V_{\theta}(x,t)$ of a model as follows. Suppose we have an input image $x$, and the diffusion step of the API is $t$. The variation API randomly adds $t$-step Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to the image and denoises it using the DDIM sampling process $\phi_{\theta}(x_{t},t)$ as described in Equation 1, returning the reconstructed image $V_{\theta}(x,t)=\hat{x}$. The details are as follows:
$x_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\quad\hat{x}=\Phi_{\theta}(x_{t},0)=\phi_{\theta}(\cdots\phi_{\theta}(\phi_{\theta}(x_{t},t),t-1),0)\,.$ (2)
This definition aligns with the image-to-image generation method of the diffusion model, making access to this API practical in many setups [27, 37, 49]. Some APIs provide the user with a choice of $t$, while others do not and use the default parameter. We will discuss the influence of different diffusion steps in Section 5.5, showing that the attack performances are relatively stable and not sensitive to the selection of $t$. We also note that for the target model, we can substitute $\phi_{\theta}(x_{t},0)$ with other sampling methods, such as the Euler-Maruyama Method [29] or Variational Diffusion Models [18].
@@ 4.2 Algorithm
Denote $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, we assume that the denoise model is expressive enough such that for the input $x_{0}\in\mathbb{R}^{d}$ and time step $t\in\mathcal{T}$, the Jacobian matrix $\nabla_{\theta}\epsilon_{\theta}(x_{t},t)$ is full rank $(\geq d)$. This suggest that the model can adjust the predicted noise $\epsilon_{\theta}(x_{t},t)$ locally in any direction. Then for a well trained model, we would have
$\nabla_{\theta}L(\theta)=0$ $\implies\nabla_{\theta}\epsilon_{\theta}(x_{t},t)^{T}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right]=0\,,$
$\implies\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right]=0\,.$
Intuitively, this is because if the noise prediction from the neural network exhibited high bias, the network could adjust to fit the bias term, further reducing the training loss.
@@
Therefore, for images in the training set, we expect the network to provide an unbiased noise prediction. Since the noise prediction is typically inaccessible in practical applications, we use the reconstructed sample $\hat{x}$ as a proxy. Leveraging the unbiasedness of noise prediction, we demonstrate that averaging over multiple independent reconstructed samples $\hat{x}_{i}$ significantly reduces estimation error (see Theorem 1). On the other hand, for images that are not in the training set, the neural network may not provide an unbiased prediction at these points. We illustrate the intuition in Figure 2.
@@
Our algorithm classifies a sample as being in the training set if $D(x,\hat{x})$ is smaller than a threshold $\tau$, where $D(x,\hat{x})$ represents the difference between the two images. It can be calculated using traditional functions, such as the SSIM metric [48]. Alternatively, we can train a neural network as a proxy. In Section 5, we will introduce the details of $D(x,\hat{x})$ used in our experiment.
Our algorithm is outlined in Algorithm 1, and we name it ReDiffuse. The key ideas of our algorithm are illustrated in Figure 1, and we also provide some theoretical analysis in Theorem 1 to support it.
@@
     Use the variation API $V_{\theta}$ to generate the variation image $\hat{x}_{k}=V_{\theta}(x,t)$ according to Equation (2).
@@
  return  "YES" if the distance between the two images $D(x,\hat{x})$ is less than $\tau$, otherwise "NO".
@@
Theorem 1.
Suppose the DDIM model can learn a parameter $\theta$ such that, for any $x\sim D_{\text{training}}$ with dimension $d$, the prediction error $\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)$ is a random variable $X=(X_{1},X_{2},\ldots,X_{d})$ with zero expectation and finite cumulant-generating function for each coordinate [8]. Suppose the sampling interval $k$ is equal to the variation API diffusion step $t$. Let $\hat{x}$ be the average of $n$ output images of $x$ using the variation API. Then we have
@@
We note that the unbiased assumption on predicted noise can be strong in practice. For experiments, we expect the mean $\|\mu\|$ of predicted noise for trained data is smaller than the mean $\|\mu^{\prime}\|$ from unseen data. As a result, the empirical best choice of $n$ would be determined by the gap $\|\mu^{\prime}\|-\|\mu\|$.
@@ 4.3 MIA on Other Diffusion Models
$z=\text{Encoder}(x)\,,\quad z_{t}=\sqrt{\bar{\alpha}_{t}}z+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\quad\hat{z}=\Phi_{\theta}(z_{t},0)\,,\quad\hat{x}=\text{Decoder}(\hat{z})\,.$ (3)

### v4

@@ 1 Introduction
Recently, there has been a surge in the popularity of generative models, with diffusion models in particular, gaining huge attention within the AI community (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Song et al., 2020b). These models have demonstrated remarkable capabilities across various tasks, including unconditional image generation (Ho et al., 2020; Song et al., 2020a), text-to-image generation (Rombach et al., 2022; Yu et al., 2022; Nichol et al., 2021) and image-to-image generation (Saharia et al., 2022a). This surge has given rise to powerful AI art models such as DALL-E 2 (Ramesh et al., 2022), Stable Diffusion (Rombach et al., 2022), and Imagen (Saharia et al., 2022b). AI-generated art holds a promising future and is expected to have a widespread impact.
Effective training of diffusion models requires high-quality data. It is thus crucial to design an algorithm that can identify whether a specific artwork has been used during the training of a model, thereby providing protection for these artworks and detecting misuse of data. This is especially important due to the rapid growth of generative models, which has raised concerns over intellectual property (IP) rights, data privacy, and the ethical implications of training on copyrighted or proprietary content without consent. As these models are increasingly deployed across industries, detecting whether a specific piece of content was used in training can help prevent unauthorized use of artistic works, protecting creators’ copyrights and ownership rights. This is a classic problem in the field of machine learning, first introduced by Shokri et al. (2017) and named “membership inference attack”.
A series of studies have been conducted on membership inference attacks against diffusion models. Hu & Pang (2023) was the first to examine this issue, utilizing the loss function values of diffusion models to determine whether an image is in the training set. Duan et al. (2023) and Kong et al. (2023) extended this work by relaxing the assumptions about model access requirements.
@@
    We evaluate our method using DDIM (Song et al., 2020a) and Stable Diffusion (Rombach et al., 2022) models on classical datasets, including CIFAR10/100 (Krizhevsky et al., 2009), STL10-Unlabeled (Coates et al., 2011), LAION-5B (Schuhmann et al., 2022), etc. Our method outperforms the previous methods.
@@
    We extend both existing algorithms and our own algorithm to the Diffusion Transformer (Peebles & Xie, 2023) architecture, implementing the membership inference attack within this model framework for the first time. Experimental results demonstrate that our algorithm is consistently effective.
@@ 2 Related Works
The diffusion model, initially proposed by Sohl-Dickstein et al. (2015), has achieved remarkable results in producing high-quality samples across a variety of domains. This ranges from image generation (Song & Ermon, 2019; Song et al., 2020b; Dhariwal & Nichol, 2021), audio synthesis (Popov et al., 2021; Kong et al., 2020; Huang et al., 2022), and video generation (Ho et al., 2022a, b; Wu et al., 2023), etc. Among existing diffusion models, the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) is one of the most frequently adopted. This approach introduces a dual-phase process for image generation: initially, a forward process gradually transforms training data into pure noise, followed by a reverse process that meticulously reconstructs the original data from this noise. Building on this model, there have been numerous follow-up studies, such as Stable Diffusion (Rombach et al., 2022), which compresses images into a latent space and generates images based on text, and the Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020a), which removes Gaussian randomness to accelerate the sampling generation process. These advancements demonstrate the versatility and potential of diffusion models.
@@
In the era of big data, preserving data privacy is paramount. The training of diffusion models may involve sensitive datasets like artists’ artworks, which are protected by copyright laws. Membership inference attacks, initially introduced by Shokri et al. (2017), serve as an effective means to detect potential misuse of data without proper authorization. Its objective is to ascertain whether a particular data sample participated in the training phase of a target model. This approach is instrumental in probing privacy breaches and identifying illicit data utilization. Researchers primarily focus on membership inference attacks for classification models (Salem et al., 2018; Yeom et al., 2018; Long et al., 2018; Li & Zhang, 2021), embedding models (Song & Raghunathan, 2020; Duddu et al., 2020; Mahloujifar et al., 2021), and generative models (Hayes et al., 2017; Hilprecht et al., 2019; Chen et al., 2020).
In the domain of membership inference attacks against diffusion models, Wu et al. (2022); Hu & Pang (2023) use a white-box approach, which assumes access to the entire diffusion model and utilizes loss and likelihood to determine whether a sample is in the training set. Duan et al. (2023); Kong et al. (2023); Tang et al. (2023) have relaxed these requirements, eliminating the need for the entire model. They leverage the insight that samples within the training set yield more accurate noise predictions, thereby achieving high accuracy in membership inference attack tasks. However, they also require the outputs of the U-net, as it is necessary to obtain the noise predictions of intermediate steps. Recently, Pang & Wang (2023) proposed a black-box membership inference attack method against diffusion models. Their method identifies whether a specific image is in a finetuning dataset of 100 images by calculating the difference between the generated image and the target image, using the corresponding prompt as input. Their approach leverages the model’s tendency to memorize finetuning images and generate similar outputs. In contrast, we focus on detecting whether an image is in the pretraining dataset, where a pre-trained model often produces diverse outputs for the same prompt, making detection more challenging.
@@ 3 Preliminary
In this section, we begin by introducing the notations used for several popular diffusion models. We first introduce the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020). Then, we extend to the Denoising Diffusion Implicit Model (DDIM) (Song et al., 2020a) and Stable Diffusion (Rombach et al., 2022), which are variants of DDPM used to accelerate image generation or generate images grounded in text descriptions. Lastly, we discuss Diffusion Transformer (Peebles & Xie, 2023), a model that replaces the U-net architecture with a transformer and achieves higher-quality image generation.
@@
$x_{t-1}=\phi_{\theta}(x_{t},t)$ $=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)$
$+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x_{t},t)\,,$
@@
Stable Diffusion leverages a variational autoencoder (VAE) (Kingma & Welling, 2013) to encode images into a latent space and perform diffusion in this compressed space. The model uses a text encoder to guide the diffusion process, enabling text-to-image generation:
@@
Diffusion Transformer leverages the Vision Transformer (Dosovitskiy, 2020) structure to replace the U-net architecture traditionally used in diffusion models for noise prediction. Its training and sampling methods remain consistent with DDIM, with the only difference being the replacement of noise prediction network $\epsilon_{\theta}$ with $\epsilon_{\tilde{\theta}}$, where $\tilde{\theta}$ represents a Vision Transformer-based architecture. This approach further enhances the generation quality and ensured that the model possesses good scalability properties.
@@ 4.1 The variation API for Diffusion Models
Most previous works on membership inference attacks against diffusion models aim to prevent data leakage and hence rely on thresholding the model’s training loss. For instance, Hu & Pang (2023) involves a direct comparison of image losses, while Duan et al. (2023); Kong et al. (2023) evaluates the accuracy of the model’s noise prediction at initial or intermediate steps. However, the required access to the model’s internal U-net structure prevents applications from copyright protection because most servers typically provide only black-box API access.
@@
Definition 4.1 (The variation API).
We define the variation API $V_{\theta}(x,t)$ of a model as follows. Suppose we have an input image $x$, and the diffusion step of the API is $t$. The variation API randomly adds $t$-step Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to the image and denoises it using the DDIM sampling process $\phi_{\theta}(x_{t},t)$, returning the reconstructed image $V_{\theta}(x,t)$. The details are as follows:
$x_{t}$ $=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,$
$V_{\theta}(x,t)$ $=\Phi_{\theta}(x_{t},0)=\phi_{\theta}(\cdots\phi_{\theta}(\phi_{\theta}(x_{t},t),t-1),0)\,.$
This definition aligns with the image-to-image generation method of the diffusion model, making access to this API practical in many setups (Lugmayr et al., 2022; Saharia et al., 2022a; Wu & De la Torre, 2023). Some APIs provide the user with a choice of $t$, while others do not and use the default parameter. We will discuss the influence of different diffusion steps in Section 5.5, showing that the attack performances are relatively stable and not sensitive to the selection of $t$. We also note that for the target model, we can substitute $\phi_{\theta}(x_{t},0)$ with other sampling methods, such as the Euler-Maruyama Method (Mao, 2015) or Variational Diffusion Models (Kingma et al., 2021).
@@ 4.2 Algorithm
Denote $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, we assume that the denoise model is expressive enough(the neural network’s dimensionality significantly exceeds that of the image data) such that for the input $x_{0}\in\mathbb{R}^{d}$ and time step $t\in\mathcal{T}$, the Jacobian matrix $\nabla_{\theta}\epsilon_{\theta}(x_{t},t)$ is full rank $(\geq d)$. This suggest that the model can adjust the predicted noise $\epsilon_{\theta}(x_{t},t)$ locally in any direction. Then for a well trained model, we would have $\nabla_{\theta}L(\theta)=0$. This implies
$\implies\nabla_{\theta}\epsilon_{\theta}(x_{t},t)^{T}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(x_{t},t\right)\right]=0\,,$
$\implies\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(x_{t},t\right)\right]=0\,.$
This is intuitive because if the noise prediction from the neural network exhibits high bias, the network can adjust to fit the bias term, further reducing the training loss.
Therefore, for images in the training set, we expect the network to provide an unbiased noise prediction. Since the noise prediction is typically inaccessible in practical applications, we use the reconstructed sample $\hat{x}$ as a proxy. By leveraging the unbiasedness of noise prediction, we show that averaging over multiple independent reconstructed samples $\hat{x}_{i}$ can significantly reduce estimation error (see Theorem 4.2). However, for images not in the training set, the neural network may not provide an unbiased prediction at these points. The intuition is illustrated in Figure 2.
@@
Table 1: Comparison of different methods on four datasets for the DDIM model. Previous methods require access to the U-Net, whereas our methods do not. We use AUC, ASR, and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
Method CIFAR10 CIFAR100 STL10
Algorithm U-Net AUC ASR TP AUC ASR TP AUC ASR TP
Loss (Matsumoto et al., 2023) $\Box$ 0.88 0.82 14.2 0.92 0.84 20.9 0.89 0.82 15.6
SecMI (Duan et al., 2023) $\Box$ 0.95 0.90 40.7 0.96 0.90 44.9 0.94 0.88 26.9
PIA (Kong et al., 2023) $\Box$ 0.95 0.89 48.7 0.96 0.90 47.0 0.94 0.87 29.8
PIAN (Kong et al., 2023) $\Box$ 0.95 0.89 50.4 0.91 0.85 39.2 0.92 0.86 28.5
ReDiffuse $\blacksquare$ 0.96 0.91 40.7 0.98 0.93 48.2 0.96 0.90 31.9
@@
Our algorithm classifies a sample as being in the training set if $D(x,\hat{x})$ is smaller than a threshold $\tau$, where $D(x,\hat{x})$ represents the difference between the two images. It can be calculated using traditional functions, such as the SSIM metric (Wang et al., 2004). Alternatively, we can train a neural network as a proxy. In Section 5, we will introduce the details of $D(x,\hat{x})$ used in our experiment.
Our algorithm is outlined in Algorithm 1, and we name it ReDiffuse. The key ideas of our algorithm are illustrated in Figure 1, and we also provide some theoretical analysis in Theorem 4.2 to support it.
@@
     Use the variation API $V_{\theta}$ to generate the variation image $\hat{x}_{k}=V_{\theta}(x,t)$ according to Definition 4.1.
@@
  return ”YES” if the distance between the two images $D(x,\hat{x})$ is less than $\tau$, otherwise ”NO”.
@@
Theorem 4.2.
Suppose the DDIM model can learn a parameter $\theta$ such that, for any $x\sim D_{\text{training}}$ with dimension $d$, the prediction error $\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)$ is a random variable $X=(X_{1},X_{2},\ldots,X_{d})$ with zero expectation and finite cumulant-generating function for each coordinate (Durret, 2010). Suppose the sampling interval $k$ is equal to the variation API diffusion step $t$. Let $\hat{x}$ be the average of $n$ output images of $x$ using the variation API. Then we have
@@ 4.3 MIA on Other Diffusion Models
$z$ $=\text{Encoder}(x)\,,\quad z_{t}=\sqrt{\bar{\alpha}_{t}}z+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,$
$\hat{z}$ $=\Phi_{\theta}(z_{t},0)\,,\quad\hat{x}=\text{Decoder}(\hat{z})\,.$